# Eigenvalue bounds for Schrödinger operators with complex potentials on compact manifolds

### v1

@@ 1. We will use the following bound on the eigenvalues $a_{j}(k)$ of $\chi P_{k}\chi$,
Moreover, there are at most $j_{0}$ eigenvalues $a_{1}(k),\ldots,a_{j_{0}}(k)$ in the interval $[a_{j_{0}}(k),a_{0}(k)]$. If we rescale $\tilde{a}_{j}(k):=a_{j}(k)/a_{0}(k)$, then
@@
Let $20<\epsilon<C_{1}/$ 2, so that $I_{\epsilon}:=[1-2\epsilon,1]$ is a subinterval of $[\tilde{a}_{j_{0}}(k),1]$ and thus contains at most $j_{0}$ distinct points of the sequence $\{\tilde{a}_{j}(k)\}_{j=0}^{d_{k}}$. Let
@@ 3.2. Optimality on Zoll manifolds
Note that, due to the additional assumption $k^{2\sigma(q)}\rho\gg 1$ in the Zoll case, the eigenvalues $\lambda_{k,j}^{2}$ of $-\Delta_{g}$ corresponding to the $k$-th cluster are contained inside the disks $\tilde{D}_{k}$ defined in (12), and these disks are mutually disjoint. The situation is similar to that on the sphere, the only difference being that a single eigenvalue is now replaced by a whole cluster of eigenvalues. For this reason, we use a different splitting $K(z)=K_{\rm r}(z)+K_{\rm nr}(z)$, namely
@@
where $E_{k,j}$ is the orthogonal projection onto the eigenfunction of $-\Delta_{g}$ corresponding to $\lambda_{k,j}^{2}$. As a consequence of (33), we have an analogous bound to (78),
@@
for $\kappa\gg 1$, or equivalently, $k\gg 1$. Since $1/q=1/p-1/p^{\prime}$, the exponent of $\kappa$ equals $n/(2q)-1/2$, which is $\sigma(q)$ for $q\leq(n+1)/2$. In the case $q\geq(n+1)/2$, we appeal to [22, (5.1.13)], which for $\kappa\gg 1$ gives the lower bound
@@
The exponent of $\kappa$ equals $(n-1)/(4q)$, which is $\sigma(q)$ for $q\geq(n+1)/2$. By a $TT^{*}$ argument, the combination of (143), (144) yields
@@
if $k\gg 1$ (we recall that $\sigma(q)>0$ for $q<\infty$). Since $K_{\rm r}(w)$ is a normal operator, the analogue of (112), with $(k+\alpha)^{2}$ in place of $k(k+n-1)$, holds, and thus (c.f. (117))
@@ 4. Scaling considerations
We assume $V\in L^{q}(LM)$ for some $q\in(n/2,\infty)$; the case $q=n/2$ can be handled similarly (under an additional smallness condition), but we will not consider it here. Theorem 1.1, and more specifically (3), implies that the spectrum of $-\Delta+V$ is contained in the set
@@ 2. If $q\in(n/2,(n+1)/2)$ and
where we used the embedding (156) to identify the $L^{q}$ norm of $V$ on $L\mathbb{T}^{n}$ with that on ${\mathbb{R}}^{n}$. Since $02\sigma(q)-n/q<$ 0 for all $q\in(n/2,\infty)$, we may choose $\epsilon$ so small (depending on $q$) that $02(1+\epsilon)\sigma(q)-\frac{n}{q}<$ 0. Hence, if there exists $L_{0}>1$ such that $z_{L}\in A_{L}$ for $L\geq L_{0}$, then
@@
Since $01-\frac{n}{q}+2\sigma(q)=$ 0 for $q\leq(n+1)/2$ and $R$ is arbitrarily large, the claim follows. ∎

### v2

@@ 1. We will use the following bound on the eigenvalues $a_{j}(k)$ of $\chi P_{k}\chi$,
Moreover, there are at most $j_{0}$ eigenvalues $a_{1}(k),\ldots,a_{j_{0}}(k)$ in the interval $a_{j_{0}}(k),a_{0}(k)]$. If we rescale $\tilde{a}_{j}(k):=a_{j}(k)/a_{0}(k)$, then
@@
Let $0<\epsilon<C_{1}/2$, so that $I_{\epsilon}:=[1-2\epsilon,1]$ is a subinterval of $\tilde{a}_{j_{0}}(k),1]$ and thus contains at most $j_{0}$ distinct points of the sequence $\{\tilde{a}_{j}(k)\}_{j=0}^{d_{k}}$. Let
@@
(121) $\max_{w\in\gamma}\|\mathcal{A}_{\rm r}(w,\kappa_{0})^{-1}(\mathcal{A}(w,\kappa_{0})-\mathcal{A}_{\rm r}(w,\kappa_{0}))\|$
(122) $\leq|\kappa_{0}|\|\mathcal{A}_{\rm r}(w,\kappa_{0})^{-1}\|\max_{w\in\gamma}\|K_{\rm nr}\|$
@@ 3.2. Optimality on Zoll manifolds
Note that, due to the additional assumption $k^{2\sigma(q)}\rho\gg 1$ in the Zoll case, the eigenvalues $\lambda_{k,j}^{2}$ of $\Delta_{g}$ corresponding to the $k$-th cluster are contained inside the disks $\tilde{D}_{k}$ defined in (12), and these disks are mutually disjoint. The situation is similar to that on the sphere, the only difference being that a single eigenvalue is now replaced by a whole cluster of eigenvalues. For this reason, we use a different splitting $K(z)=K_{\rm r}(z)+K_{\rm nr}(z)$, namely
@@
where $E_{k,j}$ is the orthogonal projection onto the eigenfunction of $\Delta_{g}$ corresponding to $\lambda_{k,j}^{2}$. As a consequence of (33), we have an analogous bound to (78),
@@
for $\kappa\gg 1$, or equivalently, $k\gg 1$. Since $1/q=1/p-1/p^{\prime}$, the exponent of $\kappa$ equals n/(2q)1/2n/(2q)-1/2, which is $\sigma(q)$ for $q\leq(n+1)/2$. In the case $q\geq(n+1)/2$, we appeal to [22, (5.1.13)], which for $\kappa\gg 1$ gives the lower bound
@@
The exponent of $\kappa$ equals (n1)/(4q)(n-1)/(4q), which is $\sigma(q)$ for $q\geq(n+1)/2$. By a $TT^{*}$ argument, the combination of (143), (144) yields
@@
if $k\gg 1$ (we recall that $\sigma(q)>0$ for $q<\infty$). Since $K_{\rm r}(w)$ is a normal operator, the analogue of (112), with $(k+\alpha)^{2}$ in place of k(k+n1)k(k+n-1), holds, and thus (c.f. (117))
@@ 4. Scaling considerations
We assume $V\in L^{q}(LM)$ for some $q\in(n/2,\infty)$; the case $q=n/2$ can be handled similarly (under an additional smallness condition), but we will not consider it here. Theorem 1.1, and more specifically (3), implies that the spectrum of $\Delta+V$ is contained in the set
@@ 2. If $q\in(n/2,(n+1)/2)$ and
where we used the embedding (156) to identify the $L^{q}$ norm of $V$ on $L\mathbb{T}^{n}$ with that on ${\mathbb{R}}^{n}$. Since $2\sigma(q)-n/q<0$ for all $q\in(n/2,\infty)$, we may choose $\epsilon$ so small (depending on $q$) that $2(1+\epsilon)\sigma(q)-\frac{n}{q}<0$. Hence, if there exists $L_{0}>1$ such that $z_{L}\in A_{L}$ for $L\geq L_{0}$, then
@@
Since $1-\frac{n}{q}+2\sigma(q)=0$ for $q\leq(n+1)/2$ and $R$ is arbitrarily large, the claim follows. ∎